Mingenback Company has 560,000 shares of $10 par value common stock outstanding. During the year Mingenback declared a 5% stock dividend when the market price of the stock was $48 per share. Two months later Mingenback declared a $.60 per share cash dividend. As a result of the dividends declared during the year, retained earnings decreased by:
Answer #1

Solution:

Stock Dividend = 560,000 *5% * $48 = $1,344,000

Cash Dividend = 560000*105% * $0.60 = $352,800

Hence, retained earnings will decrease by = $1,344,000 + $352,800 = $1,696,800